OpenStudy (anonymous):

Granite, a stone found in New Hampshire and Vermont, is 0.8% water. About how many pounds of water are there in 3,000 pounds of granite?

OpenStudy (anonymous):

Look up multiplication

OpenStudy (anonymous):

shut up feather

OpenStudy (anonymous):

Can you multiply together $$\frac{8}{10}\times3000$$

OpenStudy (anonymous):

can yo momma multiply feather???

OpenStudy (anonymous):

sorry, that was wrong

OpenStudy (anonymous):

actually hyun has it right

OpenStudy (anonymous):

actually, no. 8/10 = 80%, not 0.8! A percent is a part of 100, so: 0.8% = 0.8/100 = 0.008 So, for each pound of granite, 0.008 pounds are water, following the logic: 3000 pounds * 0.008 = your answer ;)